OpenStudy (andijo76):

listed below are the budgets ( in millions of dollars) and the gross receipts ( in millions of dollars for the randomly selected movies construct a scatterplot find the value. of the linear correlation coefficient r and find the P value using a=0.05 is there sufficient evidence to conclude that there is a linear correlation between budgets and gross receipts? do the results change if the actual budgets listed are $63,000.000 $87,000.000 $49,000.000 and so on i have the scatterplot figured out i will show what i have so far

OpenStudy (andijo76):

budget(x) 63,87,49,31,193,96,91 gross(y) 61,68,45,62,536,142,48 budget (x) =63,87,49,31,193,96,91=610 gross(y)= 61,68,45,62,536,142,48=962 x*y= 3843,5916,2205,1922,103448,13632,4368=130966 x^2= 3962,7569,2401,961,37249,9216,8281=70608 y^2= 3721,4624,2025,3844,287296,20164,2304=323978 \[x=\frac{ \Sigma x }{ n }=\frac{ 610 }{ 7 }=87,y=\frac{ \Sigma y }{ n }=\frac{ 962 }{ 7 }\]=137  now i am lost